R 2R b. od E w 2 € 3R 4R 80 mA ОА OB. 10 mA Ос OD. 100 mA 75 mA 60 mA OE 50 mA OF 130 mA OG OH. 120 mA

For the electrical grid R = 1000 Ohm and E = 250V calculate the current in the ac line
Answer #1

Answer:- E) 60 uA

Here usinh nodal analysis at point c we will have get the relation as,

(Vc- E)/R + Vc(1/4R+1/3R) + (Vc -2E)/2R = 0

Vc(1/1000 +1/4000+1/3000+1/2000) = 250/1000 +500/2000

Therefore,

Vc = 240 V

Hence the current in line ac will be,

Iac = 240/4000 = 60 mA

Hence, option E) 60mA is correct here.